# JWST Discovery of a High-Redshift Tidal Disruption Event Candidate in COSMOS-Web

### v1

@@ 6 Conclusion
  • We develop criteria for identifying candidate $z>4$ TDEs given their NIRCam colors, comparing to contaminant stars, galaxies, AGN, and supernovae. For regions of potential contamination, we pose methods for distinguishing TDEs from these contaminants. This methodology can be applied to any deep field survey, and will yield a wealth of high-redshift TDEs from the Roman High Latitude Wide Area Survey.
  • We apply this selection method to the COSMOS-Web catalog, and identify 117 sources among $>700,000$ that pass our cuts. We systematically check these sources for lack of variability using the COSMOS2020 catalog, which compiles all previous imaging of the COSMOS field, and find one source which is a confirmed transient. We call this TDE candidate HZTDE-1.
  • We model HZTDE-1 as a possible supernova, both through SED modeling and host galaxy association. We find that HZTDE-1 is $>5$ directional light radii from the nearest galaxy, making an association extremely unlikely. We find that while the SED cannot be explained by a Type I supernova, a Type II (particularly IIn) at $z\sim 2$ can explain the SED mildly well, except for the farthest infrared emission. However, it is still possible that this transient is a “hostless” luminous or superluminous Type IIn supernova.
@@
Follow-up observations of HZTDE-1, both for photometric fading and for spectroscopic information, are crucial for understanding its nature and confirming its redshift. If HZTDE-1 is a TDE, it could confirm an enhancement of the volumetric rate of TDEs in the high-redshift universe. As TDEs offer a probe of central SMBH mass independently from the virial methods that use AGN emission, a population of TDEs at high redshifts could offer insight into the nature of the SMBH mass function as a function of cosmic time. This, in turn, could place additional constraints on the methods by which SMBHs seed and grow.
@@ Appendix A Verification of optical nondetection
Because of the small angular distance between HZTDE-1 and the nearby galaxy, we wanted to verify that HZTDE-1 was not obscured by the galaxy in the lower-resolution UltraVISTA imaging. We use aperture photometry to test this as follows. First, in pixel space, we calculate the location of the centroid of the galaxy in each filter by fitting a 2D gaussian to the data in the photutils Python package (Bradley et al., 2025). Filter by filter, we find the angular separation between the galaxy centroid and the PSF-modeled coordinates of HZTDE-1 which we call $d$. Starting with the location of HZTDE-1, we place a series of apertures, each with a radius of $2$ pixels, at an angular distance $d$ from the galaxy, each offset from each other by a rotation angle of $20^{\circ}$, as shown in Figure 12. We calculate both the mean flux and the total flux within each aperture in UVISTA and find no increase in the flux at the location of HZTDE-1 compared to background in any filter. We then perform the same procedure using JWST imaging, with apertures of radius 3 pixels, as seen in Figure 12. As expected, we find a ×$\times 1000$ increase in flux relative to background in NIRCam, as seen in Figure 14.

### v4

@@ 6 Conclusion
  • We develop criteria for identifying candidate $z>4$ TDEs given their NIRCam colors, comparing to contaminant stars, galaxies, AGN, and supernovae. For regions of potential contamination, we pose methods for distinguishing TDEs from these contaminants. This methodology can be applied to any deep field survey, and will yield a wealth of high-redshift TDEs from the Roman High Latitude Wide Area Survey. This methodology could also be used to identify a range of other hostless transients in upcoming surveys.
  • We apply this selection method to the COSMOS-Web catalog, and identify 117 sources among >700,000>700,000 that pass our cuts. We systematically check these sources for lack of variability using the COSMOS2020 catalog, which compiles all previous imaging of the COSMOS field, and find one source which is a confirmed transient. We call this TDE candidate HZTDE-1.
  • We model HZTDE-1 as a possible supernova, both through SED modeling and host galaxy association. We find that HZTDE-1 is >5>5 directional light radii from the nearest galaxy, making an association extremely unlikely. We find that while the SED cannot be explained by a Type I supernova, a Type II (particularly IIn) at $z\sim 2$ can explain the SED mildly well, except for the farthest infrared emission. However, it is still possible that this transient is a “hostless” luminous or superluminous Type IIn supernova.
@@
Follow-up observations of HZTDE-1, for photometric fading and for spectroscopic information, are crucial for understanding its nature and confirming its redshift. If HZTDE-1 is a TDE, it could confirm an enhancement of the volumetric rate of TDEs in the high-redshift universe. As TDEs offer a probe of central SMBH mass independently from the virial methods that use AGN emission, a population of TDEs at high redshifts could offer insight into the nature of the SMBH mass function as a function of cosmic time. This, in turn, could place additional constraints on the methods by which SMBHs seed and grow.
@@ Appendix A Verification of optical nondetection
Because of the small angular distance between HZTDE-1 and the nearby galaxy, we wanted to verify that HZTDE-1 was not obscured by the galaxy in the lower-resolution UltraVISTA imaging. We use aperture photometry to test this as follows. First, in pixel space, we calculate the location of the centroid of the galaxy in each filter by fitting a 2D gaussian to the data in the photutils Python package (Bradley et al., 2025). Filter by filter, we find the angular separation between the galaxy centroid and the PSF-modeled coordinates of HZTDE-1 which we call $d$. Starting with the location of HZTDE-1, we place a series of apertures, each with a radius of $2$ pixels, at an angular distance $d$ from the galaxy, each offset from each other by a rotation angle of $20^{\circ}$, as shown in Figure 13. We calculate both the mean flux and the total flux within each aperture in UVISTA and find no increase in the flux at the location of HZTDE-1 compared to background in any filter. We then perform the same procedure using JWST imaging, with apertures of radius 3 pixels, as seen in Figure 13. As expected, we find a $\times 1000$ increase in flux relative to background in NIRCam, as seen in Figure 15.